OpenStudy (anonymous):

Help asap teacher now 2x+3 to third power = 125

OpenStudy (anonymous):

What is to the third power? 3 or 2x?

OpenStudy (anonymous):

If it is 3 to the third power its quite simple: We know 3 to the third power is 9, so the equation would be: 2x+9=125 9 moves to the other side to subtract 125 2x=125-9 Which is equivalent to 2x=116 Divide both sides by 2 The answer is x=58 after division Hope I helped!

OpenStudy (anonymous):

Thank you so much too bad it wasnt in time hopefully im asked the same question again thank you nevertheless